At the end of April, Kaye Company had completed Jobs 50 and 51. Job 50 is for 1,000 units, and Job 51 is for 200 units.

The following data relate to these two jobs:

On April 20, raw materials were requisitioned for production as follows: 700 units for Job 50 at $16 per unit and 1,200 units for Job 51 at $22 per unit.

During April, Kaye Company accumulated 400 hours of direct labor costs on Job 50 and 900 hours on Job 51. The total direct labor was incurred at a rate of $16 per direct labor hour for Job 50 and $20 per direct labor hour for Job 51.

The predetermined factory overhead rate is $14.00 per direct labor hour.

a. Determine the balance on the job cost sheets for Jobs 50 and 51 at the end of April.

Job 50 $
Job 51 $

b. Determine the cost per unit for Jobs 50 and 51 at the end of April. Round all answers to the nearest whole cent.

Job 50 $
Job 51 $
Answer #1
(a) Job 50 Job 51
Raw Materials $11,200 =700*16 $26,400 =1200*22
Direct Labour $6,400 =400*16 $18,000 =900*20
Factory Overhead $5,600 =400*14 $12,600 =900*14
Balance on the Job Cost Sheets $23,200 $57,000
(b)
Total Job Cost $23,200 $57,000
No of Units in Job 1000 200
Cost per unit $23.20 =23200/1000 $285 =57000/200